OpenStudy (anonymous):

Suppose that T:R^3 --> R^3 is a linear transformation and we know that T(1,1,0)=(2,1,-1), T(0,1,2)=(1,0,2), T(2,1,3)=(4,0,1). Find T(3,2,-1)

OpenStudy (anonymous):

How do I even start this question? This was in the section basis and all of a sudden it's about transformations..

OpenStudy (anonymous):

Remember that linear transformations have the property T(av) = aT(v)

OpenStudy (anonymous):

What you want to do is make (3,2-1) a linear combination of the basis vectors.

OpenStudy (anonymous):

T(av) = aT(v) T(u + v) = T(u) + T(v)

OpenStudy (anonymous):

Okay I don't really get like say T(1,1,0)=(2,1,-1), what does the numbers 1,1,0 represent and what does (2,1,-1) mean

OpenStudy (anonymous):

is T(1,1,0) just like a vector V=(a,b,c)?

OpenStudy (anonymous):

(1,1,0) is a vector in R^3

OpenStudy (anonymous):

then what does (2,1,-1) mean?

OpenStudy (anonymous):

T(1,1,0)=(2,1,-1) indicates that the linear transformation maps (1,1,0) to (2,1,-1)

OpenStudy (anonymous):

Okay so would I do addition and subtraction with T(1,1,0) T(0,1,2) and T(2,1,3) until I get T(3,2,-1)?

OpenStudy (anonymous):

Start by finding the linear combination of (1,1,0), (0,1,2) and (2,1,3) that yields (3,2,-1)

OpenStudy (anonymous):

okay so to find linear combination I would rref and find the pivots and then get the equations?

OpenStudy (anonymous):

You need to solve a system of equations. To do this you need to row reduce a matrix consisting of columns of those vectors.

OpenStudy (anonymous):

okay so I put this in a matrix sorry just want to make sure it's correct, [1 0 2 3; 1 1 1 2; 0 2 3 -1]?

OpenStudy (anonymous):

Row reduce this: \[\left[ \begin {array}{cccc} 1&0&2&3\\ 1&1&1&2 \\ 0&2&3&-1\end {array} \right] \]

OpenStudy (anonymous):

\[\left[ \begin {array}{cccc} 1&0&0&{\frac {13}{5}} \\ 0&1&0&-\frac{4}{5}\\ 0&0&1&\frac{1}{5} \end {array} \right] \]

OpenStudy (anonymous):

when you put it in matrix form in generally with v=(a,b,c) v_2=(d,e,f) etc. are the vectors always put in columns first? sometimes I do it rows..

OpenStudy (anonymous):

yeah I got that too

OpenStudy (anonymous):

This should be correct, check yourself if you like \[\frac{13}{5}\cdot (1, 1, 0) + -\frac{4}{5}\cdot(0, 1, 2) + \frac{1}{5}\cdot (2, 1, 3) = (3, 2, -1)\]

OpenStudy (anonymous):

but why is the answer (26/5,13/5,-4)?

OpenStudy (anonymous):

That's not the end of the problem. But I will explain in a moment

OpenStudy (anonymous):

So we now have a linear combination that yields the vector we want to take the image of. The vector we want to transform.

OpenStudy (anonymous):

\[T(3, 2, -1) = T\left(\frac{13}{5}\cdot (1, 1, 0) + -\frac{4}{5}\cdot(0, 1, 2) + \frac{1}{5}\cdot (2, 1, 3)\right)\]

OpenStudy (anonymous):

Are you following this so far?

OpenStudy (anonymous):

Yes I am except for the part where you got 13/5(1,1,0) + -4/5(0,1,2)..etc.

OpenStudy (anonymous):

I see how it's from the matrix but I thought the last column would be like x_4 if you know what I mean

OpenStudy (anonymous):

I will explain that at the end, lets continue

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

By linearity we know the following holds: \[\begin{eqnarray*} T(3, 2, -1) &=& T\left(\frac{13}{5}\cdot (1, 1, 0) + -\frac{4}{5}\cdot(0, 1, 2) + \frac{1}{5}\cdot (2, 1, 3)\right) \\ &=& T\left(\frac{13}{5}\cdot (1, 1, 0)\right) + T\left(-\frac{4}{5}\cdot(0, 1, 2)\right) + T\left(\frac{1}{5}\cdot (2, 1, 3)\right) \\ &=& \frac{13}{5}\cdot T\left( 1, 1, 0\right) +\left(-\frac{4}{5}\right) \cdot T\left(0, 1, 2\right) + \frac{1}{5}\cdot T\left(2, 1, 3\right) \end{eqnarray*}\]

OpenStudy (anonymous):

Can you finish the problem now?

OpenStudy (anonymous):

Yeah I think so you just take those scalars and multiply them to the corresponding vector? of the linear transformation

OpenStudy (anonymous):

and then add them up. Just one last thing would this method work for T:R^3 and R^4?

OpenStudy (anonymous):

they are in different dimensions nowT:R^3 --> R^4?

OpenStudy (anonymous):

Its the same process. And now I will explain the origin of the "magic numbers". You are solving a system of equations. You've setup a matrix A consisting of a basis. You wish to find out a linear combination of that basis yielding a certain vector so you end up with. Ax = (3, 2, -1) And you are solving for x So what you were looking at before was an augmented matrix setup to solve the system of equations.

OpenStudy (anonymous):

What! That make so much more sense so really it was Ax=(3,2,-1)?

OpenStudy (anonymous):

Where A is \[\left[ \begin {array}{cccc} 1&0&2\\ 1&1&1 \\ 0&2&3\end {array} \right]\]

OpenStudy (anonymous):

haha whoa thanks for your help it makes more sense now